OpenStudy (anonymous):

Find the equation of the line passing through (2,5) and (4,9)

OpenStudy (saifoo.khan):

Give it a try Lee.

OpenStudy (anonymous):

y=2x+1

OpenStudy (anonymous):

I got y= -7/1x - 2.5 .....

OpenStudy (anonymous):

First get the slope from the two points \[ slope(m) = \frac{y_{2} - y_{1}} {x_{2} - x_{1}} \] \[ slope(m) = \frac{9_{2} -5_{1}} {4_{2} - 2_{1}} \] Now that you have the slop use this equation \[ y - y_{1} = m(x - x_{1)} \] now just use anyone of your points \[ y - 2_{1} = m(x - 5_{1)} \] solve the equation but remember to put your slope in for M

OpenStudy (anonymous):

\[ε: y-y_0=m(x-x_0)\]\[ε:y-5= (9-5)/(4-2)\times (x-2)\]\[ε:2x-y+1=0\]or